# A Comprehensive Assessment Benchmark for Rigorously Evaluating Deep Learning Image Classifiers

### v2

@@ 3.2 Different Unknown Class Rejection Methods
| regmixup |
| Training Method | Clean | Corrupt | Adversarial | Novel | Unrecog. | Mean |
| baseline | ||||||
| noise | ||||||
@@
| baseline | .00) | |||||
| noise | .01) | |||||
@@
| AT | .21) | |||||
@@
[…]
3.3 Effects of Increasing Network Capacity
An increase in network capacity is typically accompanied by an increase in performance. Hence, in an attempt to improve on the results presented in Section 3.1 these experiments were repeated using the WideResNet34-20 architecture, that has more than 17 times the number of trainable parameters compared with the ResNet18 architecture used previously. The results for CIFAR10 are shown in Table 6. Comparing these results to those produced under equivalent conditions with ResNet18 shown in Fig. 3(c), it can be seen that, surprisingly, results for regmixup are worse for the network with higher capacity, due to a poorer ability to reject samples from unknown classes. In contrast, the overall performance improves when baseline, noise, AT, and pixmix are used as the training data augmentation methods. The improvement for pixmix is only very modest, while that of AT is the largest. The large improvement in the overall performace for AT is accompanied by a large increase in robustness to unrecognisable data. However, the results on WideResNet34-20 show that for this training method, as for all the others, there is at least one type of test data on which the model performs poorly.
Comparing the results for CIFAR100 in Table 7 and Fig. 3(f) it can be seen the the performance produced by WideResNet34-20 is superior to that produced by ResNet18 for all the different training methods, and the largest increase in performance is for regmixup. However, as for the CIFAR10 results, even with a large capacity network the robustness of the networks produced by each training method is still poor and there is still at least one type of test data for which performance is very poor.
3.4 Results for Other Data-sets
The determine how well the preceding results generalise, experiments were repeated using a more challenging data-set (TinyImageNet) and an easier one (MNIST). For MNIST different network architectures were also used. The results for TinyImageNet (Table 8) are very similar to the equivalent results produced for CIFAR100 (Fig. 3(f)), but with a small drop in overall performance for each training method. In contrast, the results for MNIST (Tables 9 and 10) differ from the results for the other data-sets and differ in details between the two architectures used. Specifically, the best overall robustness is obtained for the MNIST trained networks, as this is the least challenging data-set. However, surprisingly the best performance is not produced by pixmix, as for the other data-sets, but by training with noise.
3.5 Benchmarking Leading Methods from RobustBench
| Model | Robustness | Clean | Corrupt | Adversarial | Novel | Unrecog. | Mean |
[…]
| Bartoldson et al., (2024) | adv. () | ||||||
[…]
| Model | Robustness | Clean | Corrupt | Adversarial | Novel | Unrecog. | Mean |
[…]
| Cui et al., (2024) | adv. () |
RobustBench (Croce et al.,, 2021) is an existing benchmark for assessing the robustness of DNNs. Unlike the proposed benchmark, RobustBench evaluates performance of a model using only a single type of data: either adversarial (produced using either or constrained AA) or common corruptions. The two best performing models from each RobustBench category (for which pretrained models were available from the RobustBench repository on November 15th 2024) were evaluated using the proposed benchmarkdddAA crashed when evaluating the second best method for CIFAR100 adversarial robustness (Bai et al.,, 2024) due to the model producing NaNs. This was the case when testing using AA implemented using the torchattacks library (as has been used to adversarially attack all other methods assessed in this paper), and when using the autoattack library’s version of AA (the implementation used by RobustBench). Results for the first and third best methods are therefore presented for CIFAR100 adversarial robustness.. The results are shown in Tables 11 and 12 for the CIFAR10 and CIFAR100 data-sets, respectively. Note that these methods use the following model architectures: WideResNet18-2 (268M parameters) (Diffenderfer et al.,, 2021); ResNet18 (11M parameters) (Kireev et al.,, 2022; Modas et al.,, 2022); WideResNet94-16 (366M parameters) (Bartoldson et al.,, 2024); MeanSparse RaWideResNet70-16 (267M parameters) (Amini et al.,, 2024); WideResNet70-16 (267M parameters) (Wang et al.,, 2023; Rebuffi et al.,, 2021); WideResNet28-10 (36M parameters) (Cui et al.,, 2024).
It can be seen that the state-of-the-art RobustBench methods suffer from the same issues observed in the previous experiments. For the models trained to generalise to common corruptions the performance against adversarial attacks is poor. While the models trained to be robust to adversarial attacks are poor at rejecting unrecognisable images. Hence, overall, these models lack the robustness that conventional methods of evaluation suggest they have. On the contrary, the proposed comprehensive evaluation shows that methods of improving performance against adversarial attack, for example, fail to provide security as an attacker can use cheaply and easily generated random images to fool the classifier with a high chance of success.
@@ 4 Discussion
Despite these limitations, current practice is to test classifiers using only one (typically the clean test data), or two (clean and adversarial, or clean and novel) types of test data. While others have previously advocated using more types of test data (see Section 1.4) they have proposed using different metrics for different data-sets (Section 1.5) , which fails to adequately test true performance. This article promotes an existing metric (Zhu et al.,, 2024) that is capable of being used with different types of data to provide a consistent and comprehensive evaluation of performance.
Using the proposed benchmarkeeeNote that the proposed evaluation benchmark is likely to be imperfect, and is likely to be biased: like all assessment methods it will favour certain models over others, those that are good at processing the specific data-sets that has been chosen for performing the evaluation, but may be poorer at dealing with other data. However, the proposed benchmark is less biased than those that make no attempt to evaluate robustness to certain types of data, and (as discussed in Section 2.2) the data sets used for assessment in the proposed benchmark can easily be extended. it is found that existing methods of training DNNs, including methods that are claimed to produce state-of-the-art robustness, fail to perform as well as previous, less comprehensive, assessments would suggest. Furthermore, it appears that there are trade-offs that mean that training methods that have good performance on one type of data (i.e., improve the placement of the decision boundary in one region of feature-space) suffer from poor performance on other data (i.e., place the decision boundary less accurately elsewhere in feature-space). This means that all methods of training DNNs that have been tested here are highly susceptible to malicious attacks and, by appropriate choice of data type, can be easily fooled into producing the wrong predictions. Improved performance might be possible by combining current methods: e.g., using one model to perform unknown class rejection, and an adversarially robust model to classifying the unrejected samples. However, appropriate evaluation protocols, such as the one advocated here, will still be needed to test such ensemble methods.
@@
Without such a change in evaluation practices it will continue to be the case that the first time DNNs are rigorously assessed is after deployment in the real-world. It will therefore continue to be the case that real-world performance will be far lower than expected, and the field will continue to be rightly criticised for making exaggerated claims. Worse leaving rigorous testing until production can have dire consequences, that the research field should feel a moral duty to at least attempt to avoid. The current results show that for the MNIST data-set recent methods of training DNNs produce classifiers that are less robust than ones trained using data augmented with Gaussian noise: a technique that has been around for at least three decades. Without the advocated change in evaluation practices the field is likely to continue to fail to make advances in improving the security, reliability and trustfulness of machine learning models. Worse, lots of time, money, and CO2 (Strubell et al.,, 2020; Thompson et al.,, 2020) will continue to be expended creating methods that give a false sense of progress and security, but when tested thoroughly do not produce improved robustness.

### v3

@@ 3.2 Experiments to Evaluate Model Robustness
| Kireev et al., (2022) | ||||||||
@@
| Amini et al., (2024) |
| Model | Params. | Robustness | Clean | Corrupt | Adv. | Novel | Unrecog. | Mean |
@@
| Debenedetti et al., (2023) |
| Model | Params. | Robustness | Clean | Corrupt | Adv. | Novel | Unrecog. | Mean |
@@
| Singh et al., (2023) |
@@ 4 Discussion
Despite these limitations, current practice is to test classifiers using only one (typically the clean test data), or two (clean and adversarial, or clean and novel) types of test data. While others have previously advocated using more types of test data (see Section 1.4) they have proposed using different metrics for different data-sets (Section 1.5), which fails to adequately test true performance. This article promotes an existing metric (Zhu et al.,, 2024) that is capable of being used with different types of data to provide a consistent and comprehensive evaluation of performance. The proposed benchmark is likely to still be imperfect, and is likely to be biased: like all assessment methods it will favour certain models over others, those that are good at processing the specific data-sets that has been chosen for performing the evaluation, but may be poorer at dealing with other data. However, the proposed benchmark is less biased than those that make no attempt to evaluate robustness to certain types of data, and (as discussed in Section 2.2) the data sets used for assessment in the proposed benchmark can easily be extended.
Using the proposed benchmark it is found that existing methods of training DNNs, including methods that are claimed to produce state-of-the-art robustness, fail to perform as well as previous, less comprehensive, assessments would suggest. Furthermore, it appears that there are trade-offs that mean that training methods that have good performance on one type of data (i.e., improve the placement of the decision boundary in one region of feature-space) suffer from poor performance on other data (i.e., place the decision boundary less accurately elsewhere in feature-space). This means that all methods of training DNNs that have been tested here are highly susceptible to malicious attacks and, by appropriate choice of data type, can be fooled into producing the wrong predictions with high probability. Improved performance might be possible by combining current methods: e.g., using one model to perform unknown class rejection, and an adversarially robust model to classify the unrejected samples. However, appropriate evaluation protocols, such as the one advocated here, will still be needed to test such ensemble methods.
@@
Without such a change in evaluation practices it will continue to be the case that the first time DNNs are rigorously assessed is after deployment in the real-world. Real-world performance will be far lower than expected, and the field will continue to be rightly criticised for making exaggerated claims. Worse leaving rigorous testing until production can have dire consequences, that the research field should feel a moral duty to at least attempt to avoid. Without the advocated change in evaluation practices the field is likely to continue to fail to make advances in improving the security, reliability and trustfulness of machine learning models. Worse, lots of time, money, and CO2 (Strubell et al.,, 2020; Thompson et al.,, 2020) will continue to be expended creating methods that give a false sense of progress and security, but when tested thoroughly do not produce substantially improved robustness.
@@ Acknowledgements
Appendix A Appendix: Methods used to Train Models
A.1 Training-Time Data Augmentations
- baseline
Improved performance on the standard test data is often achieved by applying simple geometric data augmentations to the training data. Here 4-pixel random cropping and random horizontal flipping where applied to the CIFAR10, CIFAR100 and TinyIN training data, while no augmentations were applied to the MNIST data. These standard augmentations were used in combination with the additional augmentation methods listed below. Given that these baseline augmentation methods are very commonly used when training models, the results show how standardly trained networks cope under comprehensive evaluation. The results for this method of training also provides a baseline against which to judge the effectiveness of other methods that have been proposed to improve generalisation performance.
- noise
Augmenting training images with random noise has been shown to reduce over-fitting, and hence, improve generalisation (Sietsma and Dow,, 1991; Holmstrom and Koistinen,, 1992; Goodfellow et al.,, 2016), especially to common image corruptions (Rusak et al.,, 2020; Gilmer et al.,, 2019). Various methods of adding noise to the training data have been proposed. Rusak et al., (2020) augmented half the training data with Gaussian noise, and found that this technique produced performance on corrupt test data that was competitive with the more complex augmentation method, AugMix, and that for the MNIST data-set specifically was just as effective as their proposed method of Adversarial Noise Training. Lopes et al., (2019) advocated adding Gaussian noise to only a randomly chosen patch of the image, which they claimed produced better robustness to image corruptions on CIFAR10 data than applying random noise to the whole image. Gilmer et al., (2019) used an alternative method for corrupting training images with Gaussian noise in which the standard deviation of the noise applied to each image was chosen with equal probability from the range . The method implemented here is a generalisation for the method proposed by Gilmer et al., (2019). Specifically, for each image the standard deviation of the noise was chosen from a uniform distribution with range . Then each pixel intensity value was independently modified by the addition of a value chosen at random from a Gaussian distribution with mean of zero and the chosen standard deviation. The original images had pixel intensity values in the range [0,1], and after the application of the noise pixel values were clipped at 0 and 1 to remain in this range. Appropriate values of and were found via a grid search using small ConvNets. Using this method and were chosen for use with CIFAR10, and the same hyper-parameters were also used for CIFAR100 and TinyIN. For MNIST was used.
- AT
Multi-step adversarial training with Projected Gradient Descent (PGD; Madry et al.,, 2018) is a standard, and highly effective, defence against adversarial attack. As a result, this method has become a standard benchmark against which all other methods of adversarial defence are judged. This method was, therefore, also selected as a data augmentation technique to be tested. It was implemented using the torchattacks PyTorch library (Kim,, 2021). To perform adversarial training, clean images from the current batch of training examples were used to generate non-targeted adversarial examples, that then replaced the original training images (as in Madry et al.,, 2018). Following standard protocols, the maximum allowed perturbation was constrained by the -norm to be less than the perturbation budget (), and the training-time adversarial images were generated using 10 steps of PGD (denoted as PGD for brevity). Standard values used in the previous literature were used for . Specifically, was set to for performing AT with the CIFAR10, CIFAR100 and TinyIN training data, and was used for MNIST.
- pixmix
- regmixup
A.2 Training Setup
The ResNet18 (He et al.,, 2016) architecture was chosen as this is very commonly used in the literature. This model has just over 11M trainable parameters. Training was performed using a batch size of 128 for 110 epochs using the SGD optimiser with weight decay of 5e-4 and an initial learning rate of 0.1. The learning rate was step-wise decayed by a factor of 10 at the start of epochs 100 and 105. The set-up was based on that recommended by Pang et al., (2021) for performing AT on the CIFAR data-sets. Hence, these hyper-parameters may not be optimal for training with the other data augmentations used.
For each experimental condition (combination of data-set, network architecture, and training data augmentation method) the network was trained and evaluated three times, each time with a different random weight initialisation and random presentation order of training samples. The reported result for each condition is the mean accuracy over these three trials. The standard-deviation in the mean performance measured using the proposed metric is shown in the tables of results and generally shows a high degree of consistency across trials. Publicly available checkpoints created by other researchers were only evaluated once.
Open-source code written in Python using the Pytorch library (Paszke et al.,, 2019), which can be used to train networks and perform all the experiments described in this article, is available for download from: https://codeberg.org/mwspratling/RobustnessEvaluation.